OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)? Q(1, 11) R(2, 4) S(4, -4) T(9, -2)

OpenStudy (anonymous):

The general equation for a circle centered at (h,k) with radius, R, is (x-h)^2+(y-k)^2=R^2 In this case, (x-6)2+(y-1)^2=5^2 (x-6)^2+(y-1)^2=25 Now test each point with this formula and see which ones make the formula correct. Example: (1,11) (1-6)^2+(11-1)^2=25 (-5)^2+(10)^2=25 25+100=25 125=25 False, so point Q is not on the circle. Do the same for the remaining points.